# Perception Stitching: Zero-Shot Perception Encoder Transfer for Visuomotor Robot Policies

### v1

@@ Appendix C Quantitative Analysis of the Module Interface
where $g_{1}^{E}$ and $g_{2}^{E}$ are the visual encoders for $\mathbb{O}_{1,E,T}$ and $\mathbb{O}_{2,E,T}$ separately, $S_{C}(\boldsymbol{a},\boldsymbol{b})=\frac{\boldsymbol{a}\boldsymbol{b}}{\|\boldsymbol{a}\|\|\boldsymbol{b}\|}$ is the cosine similarity and $d_{L2}(p,q)=\|p-q\|$ is the L2 distance. Fig. 6 shows the cosine and L2 distances in all the experiments in the Push task. We record the distances data and the mean cosine and L2 distances across all these experiments in Table 6.
@@
Appendix D Perception Stitching with Single Camera
Push Lift Can Stack Door Open Average
This section aims to answer the question: can PeS work with policies trained with only one camera?
We train two policies with two different cameras separately. Each policy still has two visual encoders, but one visual encoder takes in a black image and the other encoder takes in the images from the single camera. For both encoders of the policy, they use the same anchor images set collected by the only camera, because we assume that each policy is trained separately and only has access to the dataset of one camera.
In the perception stitching process, we stitch the visual encoder of policy 2 to the visual encoder and action decoder of policy 1, as shown in Figure 10, and assemble a stitched policy for double camera inputs. We believe this is the only meaningful setting for the perception stitching of single-camera policies. If each policy has only one encoder and one decoder and does not have a black image encoder, we can stitch encoder 1 to decoder 2, but the stitched policy will still be applied in the environment where we train policy 1. There is no meaning to use a stitched policy when we already have a policy 1 trained to 100% of success rate in this environment.
Appendix E Perception Stitching with Multiple Cameras
Mask Zoom in Blurred Noise Fisheye Camera Position Average
This section aims to answer the question: can PeS work with policies trained with more than two cameras?
We pick the Stack task and train policies with three visual encoders that embed three different visual configurations. For the first five visual configurations that add different effects to the observations, policies are all trained with the in-hand view, close-up view, and side view. Policy 1 has some effect (e.g. Gaussian noise) only in vision 3, and policy 2 has the same effect in vision 1 and vision 2. Then we stitch the encoder 3 of policy 2 to the rest parts of policy 1 and test the stitched policy with three normal cameras. For the last experiment that involves different camera positions, policy 1 is trained with the in-hand view, close-up view, and front view, and policy 2 is trained with close-up view, front view, and side view. Then we stitch the side view encoder of policy 2 to the in-hand view, front view encoders, and decoder of policy 1. The stitched policy is tested with agent view, front view, and side view. This triple camera perception stitching process is presented in Figure 11.
Appendix F Influence of Using Decoder 1 and Decoder 2
@@
Appendix G Anchors from failure trajectories
@@
Appendix H Highlight Attention Regions with Grad-CAM
We modify the Gradient-weighted Class Activation Mapping (Grad-CAM) [74] approach to highlight the regions that the policies pay attention to. Grad-CAM is widely applied to neural classification models with convolutional layers. To adapt the Grad-CAM from image classification to robot learning, we replace the before-softmax score $y^{c}$ for class $c$ of the image classification networks with the log-likelihood $l(a)$ of the robot action $a$ in the training dataset. We denote the $k^{th}$ feature map activation output from the last convolutional layer as $A^{k}$. Then the backpropagated gradient of $l(a)$ with respect to $A^{k}$ is computed as $\frac{\partial l(a)}{\partial A^{k}}$. We do global average pooling of these gradients over the width (indexed by $i$) and height (indexed by $j$) dimensions of the feature map to get the neuron importance weight :
This weight $\alpha^{a}_{k}$ captures the ‘importance’ of feature map $k$ for robot action $a$. Then, the attention map $L_{\text{Grad-CAM }}^{a}$ is calculated as the weighted combination of forward activation maps followed by a ReLU [74]:
We apply ReLU because we are only interested in the features that have a positive influence on the actions. The intensity of these pixels should be increased in order to increase the log-likelihood $l(a)$ [74]. This $L_{\text{Grad-CAM }}^{a}$ is a heatmap of the same size as the convolutional feature maps $A^{k}$. We upsample it to the input image size with bilinear interpolation to get the final attention heatmap of the input image [74]. A larger value on this heatmap means this pixel contributes to a larger gradient of the log-likelihood of the robot action.

### v2

@@ Appendix C Quantitative Analysis of the Module Interface
where $g_{1}^{E}$ and $g_{2}^{E}$ are the visual encoders for $\mathbb{O}_{1,E,T}$ and $\mathbb{O}_{2,E,T}$ separately, $S_{C}(\bm{a},\bm{b})=\frac{\bm{a}\bm{b}}{\|\bm{a}\|\|\bm{b}\|}$ is the cosine similarity and $d_{L2}(p,q)=\|p-q\|$ is the L2 distance. Fig. 8 shows the cosine and L2 distances in all the experiments in the Push task. We record the distances data and the mean cosine and L2 distances across all these experiments in Table 8.
@@
Appendix D Influence of Using Decoder 1 and Decoder 2
@@
Appendix E Anchors from failure trajectories
@@
Appendix F Latent Representations Visualization
We visualize the latent representations of the Lift, Can, Stack and Door tasks. Among all the visual configuration changing experiments, we choose the camera position variation experiments. We first reduced the 256D representations to 3D with PCA (Hotelling, 1933) for visualization. Since the side view encoder of the policy 2 is stitched to the policy 1 at the position of its original front view encoder, we compare the latent representations of these two encoders.
In these visualizations, the red data points are the first $5$ steps of images in each of the $200$ games in the dataset of a certain task. The blue data points are the last $5$ steps of images, and the green data points are the $5$ steps of images in the middle of each trajectory. Therefore, we have $1000$ data points for each color which represent the starting, middle, and ending stages of a task.
The visualization results show that PeS can better align the latent space and force an approximate invariance of the latent representations. In contrast, the Devin et al. (2017) baseline without adopting the relative representation and the disentanglement regularization leads to different latent representations between the two encoders, and they have an approximately isometric transformation relationship. These visualizations support our conclusions in the paper.
Appendix G Impact of Anchor Number
To study the impact of number of anchors on transfer performance, we picked a challenging task, Stack, and reported the success rates with standard errors over 3 random seeds on all the different visual configurations in Table 11. We found that when the anchor number is too small, the latent space of the visual encoder does not have enough capacity to capture effective visual representations while maintaining an approximate invariance at the same time. Therefore, the zero-shot transfer performance drops drastically. On the other hand, when the number of anchors becomes too large, there are more redundant anchors which are similar to each other. This will make disentangling the features at the latent space with the disentanglement loss harder, which can also lead to some drop of the performance. We empirically found that an anchor number between 256 to 512 can achieve optimal performance in our tasks.
Anchor Number Masked Zoom in Blurred Noise Fisheye Position Average
1024 72.7$\pm$0.94 56.0$\pm$4.32 88.7$\pm$3.40 86.7$\pm$3.77 65.3$\pm$2.49 69.3$\pm$2.49 73.1
512 90.7$\pm$3.40 94.7$\pm$0.94 91.3$\pm$0.94 92.7$\pm$5.73 95.3$\pm$3.40 88.7$\pm$9.84 92.2
256 94.7$\pm$0.94 96.7$\pm$0.94 90.0$\pm$1.63 96.7$\pm$1.89 97.3$\pm$2.49 80.0$\pm$4.90 92.6
128 72.7$\pm$10.87 43.3$\pm$5.73 42.0$\pm$1.63 24.0$\pm$2.83 73.3$\pm$1.89 30.0$\pm$7.12 47.6
64 1.3$\pm$0.00 6.7$\pm$0.94 6.7$\pm$4.11 0.0$\pm$0.00 5.3$\pm$2.49 0.0$\pm$0.00 3.3
Table 11: Success rates of the Stack task with different anchor numbers
We also investigated the impact of the anchor number on computational efficiency. We picked the Stack task and trained the policies on a NVIDIA A6000 GPU with the batch size of $32$. As shown in Table 12, a larger anchor number will lead to a larger model size, longer anchor searching time, and longer training time for each epoch.
Anchor Number Model Size Anchors Searching Time (s) Training Time Per Epoch (s)
1024 26,835,803 1455 182
512 25,720,667 922 107
256 25,163,099 570 66
128 24,884,315 288 33
64 24,744,923 194 23
Table 12: network model sizes, anchors searching time, and training time per epoch of the Stack task with different anchor numbers
To sum up, either a too large or too small anchor number can hurt the transfer performance, and larger anchor numbers can reduce the computational efficiency. It is important to select an appropriate anchor number for each task.
Appendix H Impact of Disentanglement Loss Weight
Weights Masked Zoom in Blurred Noise Fisheye Camera Position Average
0.0 34.0$\pm$11.43 10.7$\pm$4.11 62.0$\pm$10.71 34.0$\pm$7.12 22.7$\pm$3.77 26.0$\pm$4.32 31.6
0.0002 34.7$\pm$1.89 8.0$\pm$1.63 88.0$\pm$2.83 30.0$\pm$1.63 16.0$\pm$2.83 24.7$\pm$2.49 33.6
0.002 94.7$\pm$0.94 96.7$\pm$0.94 90.0$\pm$1.63 96.7$\pm$1.89 97.3$\pm$2.49 80.0$\pm$4.90 92.6
0.02 90.7$\pm$2.49 80.7$\pm$10.87 92.0$\pm$4.90 88.0$\pm$3.27 81.3$\pm$5.25 70.0$\pm$4.32 83.8
0.2 35.3$\pm$3.40 56.0$\pm$1.63 88.7$\pm$2.49 31.3$\pm$0.05 48.0$\pm$1.63 16.0$\pm$4.32 45.9
Table 13: Success rates of the Stack task with different disentanglement loss weight
We have also studied the effect of different weights of disentanglement loss. We picked the Stack task and tested different weights on all the various visual configurations. When the weight is close to zero, the transfer performance gets close to that of the PeS (w/o disent. loss) ablation method and is significantly lower than the PeS full method. When the weight becomes too large, the optimization process leans too much to disentangling the latent features than imitating the expert behaviors. Therefore, the weaker imitation of the expert agent will also cause the performance drop. We empirically found that the optimal weight of the disentanglement loss in our tasks should be around the range of $0.002$ to $0.02$. The success rates with standard errors over $3$ random seeds are shown in Table 13.
Appendix I Impact of Replay Trajectory Deviations
One limitation of PeS is that it requires a trajectory replay to collect the anchor images. However, in real-world applications, it is usually hard to accurately replay the exact trajectories. In order to understand how the trajectory deviation errors in replay could influence the performance of PeS, we conduct additional experiments to introduce trajectory deviations with different amplitudes.
We use the Stack task in the simulation to test the effect of trajectory deviations. During the trajectory replaying, we add a random horizontal vector to every position point on the trajectory except for the gripper close or open action point. The length of the random vector to cause deviation is sampled from a uniform distribution within a certain range, and we test out the range options of 0 to1 cm, 0 to 3 cm, and 0 to 5 cm. The direction of the deviation vector is randomly sampled within the x-y plane.
Amplitude Masked Zoom in Blurred Noise Fisheye Position Lighting Average
0 cm 94.7$\pm$0.94 96.7$\pm$0.94 90.0$\pm$1.63 96.7$\pm$1.89 97.3$\pm$2.49 80.0$\pm$4.90 82.0$\pm$1.63 91.1
1 cm 72.0$\pm$5.89 86.7$\pm$4.71 77.3$\pm$6.18 88.0$\pm$1.63 93.3$\pm$0.94 74.7$\pm$7.36 78.0$\pm$4.32 81.4
3 cm 34.7$\pm$3.77 35.3$\pm$1.89 44.7$\pm$8.22 26.7$\pm$1.89 37.3$\pm$2.49 19.3$\pm$7.36 23.0$\pm$2.83 31.6
5 cm 12.0$\pm$2.83 14.7$\pm$4.99 16.7$\pm$0.94 23.3$\pm$3.40 16.7$\pm$3.40 8.3$\pm$1.89 3.3$\pm$0.94 13.6
Table 14: Success rates of the Stack task with different replay trajectory deviation amplitudes
The experiment results are shown in the table below. When the trajectory deviation is within the range of 1 cm, The average performance of PeS drops by about 10%, but it can still achieve 81.4% of average success rate, which is much higher than all the baselines that don’t require trajectory replay and all the ablation methods that require trajectory replay without deviation. When the trajectory deviation goes up to the range of 3cm, PeS can achieve an average success rate of 31.6%, which is on par with the best performing baselines RT1 (29.5%) and Cannistraci et al. 2024 (linear) (34.6%). When the trajectory deviation goes up to a very large range of 5cm, the average success rate of PeS drops to 13.6%.
In summary, although PeS currently requires trajectory replays for anchor selection, it doesn’t require very precise replay and can perform well within 1 cm of replay error range.